The density of a metal is d and specific resistance is row. A wire of length l and resistance R is to prepared . How much mass would be needed for this

Grade:12

1 Answers

ashish
6 years ago
Dear jitender
as density =mass/volume   
mass =density *volume
we have   R=row*l/a
by multiplying l in numerator and in denominator
we get R=row*l2/al => R=row *l2/V (V=volume)  and (al=volume)
so V=row*l2/R put this into formula=> mass = dR /row*l2